Find the solution of the inequality:
-1 <= 4x-1/3 < 5

Answer by checkley77(12844)

-1 <= 4x-1/3 < 5
-1+1/3<=4X<5+1/3
-2/3<=4X<16/3
(-2/3)/4<=X<(16/3)/4
-2/3*1/4<=X<16/3*1/4
-2/12<=X<16/12
-1/6<=X<4/3 ANS.